# Unifying Summary Statistic Selection for Approximate Bayesian Computation

### v3

@@ Abstract
Extracting low-dimensional summary statistics from large datasets is essential for efficient (likelihood-free) inference. We characterize different classes of summaries and demonstrate their importance for correctly analysing dimensionality reduction algorithms. We demonstrate that minimizing the expected posterior entropy (EPE) under the prior predictive distribution of the model subsumes many existing methods. They are equivalent to or are special or limiting cases of minimizing the EPE. We offer a unifying framework for obtaining informative summaries, provide concrete recommendations for practitioners, and propose a practical method to obtain high-fidelity summaries whose utility we demonstrate for both benchmark and practical examples.
@@ 1 Introduction
Empowered by advances in both scientific understanding and computing, researchers are developing ever more sophisticated simulators. For example, simulated weak lensing maps capture how dark matter affects light propagating through the universe (Merten et al., 2019; Fluri et al., 2021), coalescent simulators predict the evolution of genetic material (Nordborg, 2019), and synthetic networks shed light on political opinion formation (Sobkowicz et al., 2012), effective vaccination strategies (Yang et al., 2019), and interactions between proteins (Grassmann et al., 2024).
While simulators can generate data $y$ given parameters $\theta$, we are often interested in the inverse problem: Constraining parameters $\theta$ given data $y$. If the likelihood $g\left(y\mid\theta\right)$ is available, we can use Markov chain Monte Carlo samplers (Carpenter et al., 2017) or variational inference (Bishop, 2006, Ch. 10) to investigate the posterior $f\left(\theta\mid y\right)$. But inference is more challenging if the likelihood is intractable or costly to evaluate.
Approximate Bayesian computation (ABC) overcomes this challenge in three steps by comparing observed with simulated data (Beaumont, 2019): First, we draw many samples $\left(\theta_{i},z_{i}\right)$ from the prior predictive distribution which form the so-called reference table. Second, we evaluate the distance $d_{i}=d\left(y,z_{i}\right)$ between observed data $y$ and the $i^{\mathrm{th}}$ simulated dataset $z_{i}$. Finally, we accept $\theta_{i}$ as a sample from the approximate posterior $\tilde{f}\left(\theta\mid y\right)$ if the distance $d_{i}$ is smaller than a threshold $\epsilon$. The smaller $\epsilon$, the better the approximation. Intuitively, ABC samples parameters $\theta_{i}$ that generate data $z_{i}$ which “look like” the observed data $y$. Hereafter, $y$ and $z$ will denote observed and simulated data.
@@
Figure 1: Different methods for compressing data to informative summaries are intimately related; distinguishing between classes of summaries is essential. Panel (a) illustrates that five information-theoretic approaches (ITAs) are equivalent. They implicitly minimize the same loss (Sections 2 and 3). Approximate sufficiency (Section 4.1) seeks to achieve lossless compression, and minimizing the posterior entropy (Section 4.2) is a special case of ITAs focused on only the observed data. Maximizing Fisher information (Section 4.3) and minimizing $L^{2}$ Bayes risk (Section 4.4) are equivalent each other and ITAs in the large-sample limit. Probabilistic model selection (Section 4.6) maps onto ITAs if we treat model labels as parameters. A dashed arrow from one method to another indicates that the latter is a specialization of the former. Solid arrows indicate correspondence in the large-sample limit. Panel (b) illustrates relationships between classes of summaries. Sufficient statistics $\mathcal{S}$ are a subset of lossless statistics $\mathcal{L}$ although the former only exist if the likelihood belongs to the exponential family. The intersection of $\mathcal{L}$ and the space of summaries $\mathcal{T}$ considered by the practitioner are optimal summaries $\mathcal{O}$. Optimal summaries need not be lossless such that $\mathcal{O}\cap\mathcal{L}=\varnothing$, e.g. if $\mathcal{T}$ is restricted to parametric transformations.
@@
A plethora of methods has been developed to address this question; some are summarized in panel (a) of Fig. 1. They include methods to select informative summaries from a pool of candidates (Blum & François, 2010; Joyce & Marjoram, 2008; Nunes & Balding, 2010; Barnes et al., 2012; Blum et al., 2013) and parameterized transformations that can be optimized to learn summaries (Aeschbacher et al., 2012; Fearnhead & Prangle, 2012; Prangle et al., 2014; Jiang et al., 2017; Chan et al., 2018; Charnock et al., 2018; Chen et al., 2021; Radev et al., 2022). Loss functionals quantifying how well the compressor preserves information have been motivated by minimizing the Bayes risk (Fearnhead & Prangle, 2012; Jiang et al., 2017), model selection (Prangle et al., 2014; Raynal et al., 2023; Merten et al., 2019), and information theoretic arguments (Nunes & Balding, 2010; Chen et al., 2021; Barnes et al., 2012; Charnock et al., 2018; Radev et al., 2022).
@@
After formalizing the notion of different summaries in Section 2, we argue that all information-theoretic approaches are equivalent in Section 3. They implicitly minimize the same loss functional between the posterior $f\left(\theta\mid t\right)$ given only summaries $t$ and the posterior $f\left(\theta\mid y\right)$ given the entire dataset $y$. While these results are well established in information theory, they provide a unifying perspective of different summary extraction approaches. Minimizing the expected posterior entropy (EPE) should be the practitioner’s choice because it is easier to evaluate than the Kullback-Leibler (KL) divergence or mutual information (MI). It also has strong connections with conditional posterior density estimation (Papamakarios & Murray, 2016; Lueckmann et al., 2017). But even methods developed to address different problems (such as parameter inference or model selection) in diverse fields (such as cosmology or population genetics), have strong ties to information-theoretic approaches. For example, in Section 4 we show that maximizing the determinant of the Fisher information (Heavens et al., 2000; Charnock et al., 2018) and minimizing the $L^{2}$ Bayes risk (Fearnhead & Prangle, 2012; Jiang et al., 2017) are both equivalent to minimizing the EPE in the large-sample limit. Similarly, learning a probabilistic classifier for model selection (Prangle et al., 2014) minimizes the EPE. In Section 5, we discuss concrete steps for learning summaries by fitting conditional posterior density estimators to simulated data. To compare different methods, we devise a benchmark problem with simple likelihood but data that prove challenging for summary selection in Section 5.2. We also compare summary selection approaches on two applied examples: Inferring the mutation and recombination rates of a population genetics model (Section 5.3) and the attachment kernel for a model of growing trees (Section 5.4).
@@ 2 Background
Given data $y$ we seek to infer parameters $\theta$ of a model using summaries $t=t(y)$ that retain as much information about the posterior as possible. Summaries $t_{\text{suff}}$ with fixed and finite dimensions are Bayes sufficient if $f\left(\theta\mid t_{\text{suff}}\right)=f\left(\theta\mid y\right)$ for all $y$ and any prior $\pi\left(\theta\right)$ (Prangle, 2018). But they only exist for exponential-family likelihoods (Koopman, 1936). We have to relax the concept of sufficiency, and we call statistics $t_{\text{lossless}}$ lossless if
@@
for all data $y$ of the same sample size, and a given prior $\pi\left(\theta\right)$. While lossless statistics always exist (e.g. the identity map), they may not be useful in practice. We say that the statistics $t_{\text{opt}}$ are optimal if they minimize a non-negative loss functional $\ell$ which measures the discrepancy between the posterior given the data $y$ and the posterior given only summaries $t_{\text{opt}}$. Loss functionals include, for example, the KL divergence, Wasserstein distance, and total variation distance (Cai & Lim, 2022). In particular,
$t_{\text{opt}}=\operatorname{argmin}_{t\in\mathcal{T}}\ell\left\{f\left(\theta\mid y\right),f\left(\theta\mid t(y)\right)\right\},$ (2)
@@
where $\mathcal{T}$ is the space of summaries under consideration, and the loss functional $\ell$ is zero if and only if the statistics are lossless. Consequently, sufficient statistics are lossless, and lossless statistics are optimal, but the converse is not necessarily true. For example, $\mathcal{T}$ may be restricted to parametric transformations (Fearnhead & Prangle, 2012) or selecting at most $k$ summaries from a set of candidate statistics (Raynal et al., 2023). The relationship between different classes of summaries is illustrated in panel (b) of Fig. 1.
@@
Despite the pursuit of the holy grail of sufficient statistics, we typically have to settle for the weakest concept of optimal statistics. Even the most sophisticated method cannot extract sufficient statistics if the likelihood does not belong to the exponential family (Koopman, 1936). Similarly, unless the family of summaries $\mathcal{T}$ is rich enough, lossless compression is not achievable. Further, even if $\mathcal{T}$ is rich enough, one cannot in general verify that Eq. 1 holds for all $\theta$ and $y$ given a finite computational budget.
@@
While models with exponential-family likelihoods are theoretically appealing, they may not be sufficiently expressive or intuitive to address real-world problems. Domain knowledge can aid in the development of models that capture salient features of the data, including protein interaction networks (Grassmann et al., 2024), cosmology (Charnock et al., 2018), and population-genetics (Nordborg, 2019). But these models often do not have sufficient statistics or even tractable likelihoods, and we need to resort to possibly lossy compression and likelihood-free inference.
@@ 3 Minimizing the expected posterior entropy
We propose choosing summaries that minimize the expected posterior entropy (EPE), where the expectation is taken with respect to the prior predictive distribution of the model, similar to the evaluation of the Fisher information (Bishop, 2006, Ch. 6). This approach is conceptually simple and computationally tractable, has a strong connection with recent inference techniques based on conditional density estimation (Papamakarios & Murray, 2016; Lueckmann et al., 2017; Radev et al., 2022), and is equivalent to other information-theoretic approaches, as illustrated in panel (a) of Fig. 1.
@@
Taking the expectation with respect to the data under the model yields the EPE
@@
where the Jacobian has been absorbed by the joint density $p\left(t,\theta\right)$. With a slight abuse of notation, we use $p\left(\cdot\right)$ for both the marginal likelihood and joint distribution where the distinction is unambiguous. Given a conditional density estimator $\hat{f}\left(\theta\mid t\right)$ that seeks to approximate the posterior, we can construct a Monte Carlo estimate of the EPE
^$\hat{\mathcal{H}}{}=-m^{-1}\sum_{i=1}^{m}\log\hat{f}\left(\theta_{i},t(z_{i})\right),$ (4)
where $\theta_{i}$ and $z_{i}$ are joint samples from $p\left(\theta,z\right)$, and $m$ is the number of samples. This estimate is the widely used loss function for learning the posterior from simulated data (Papamakarios & Murray, 2016; Lueckmann et al., 2017; Radev et al., 2022), where $m$ is the size of the mini-batch, i.e. a subset of the data used to train the model.
We consider three well-established connections to other information-theoretic approaches (Bishop, 2006, Ch. 1) although with a specific focus on the selection of summaries for ABC. First, we evaluate the difference between the prior entropy and EPE
@@
by the law of total probability. The inner integral of Eq. 5 is the KL divergence from the prior to the posterior $D_{\mathrm{KL}}\left(f\left(\theta\mid t\right)\;\|\;\pi\left(\theta\right)\right)$, sometimes called surprise because it measures the degree to which an observer updates their belief in light of new data (Itti & Baldi, 2009). Minimizing the EPE thus maximizes our expected surprise from observing the summaries because the prior entropy does not depend on the choice of summaries.
@@
As the MI is non-negative, the EPE is not larger than the prior entropy, i.e. we reduce uncertainty on average. Minimizing the EPE is equivalent to maximizing the MI which has been proposed in the context of subset selection (Barnes et al., 2012) and neural summaries (Chen et al., 2021). However, estimating MI is difficult in high dimensions (Jeffrey et al., 2020), making the approach computationally challenging.
@@
and we can identify the inner integral as the KL divergence from $f\left(\theta\mid t\right)$ to $f\left(\theta\mid z\right)$ (see App. A). The difference of expected entropies is thus equal to the expected KL divergence between the posteriors
@@
which Chan et al. (2018) used to infer recombination hotspots in population genetics and Radev et al. (2022) targeted for amortized Bayesian inference. Minimizing the EPE is equivalent to minimizing the expected KL divergence because the posterior entropy given the complete dataset does not depend on the summaries. The KL divergence is non-negative which allows us to draw two conclusions. First, the EPE given only summaries $t\left(z\right)$ is greater than or equal to the EPE given the full dataset $z$, i.e. we lose information in expectation by conditioning on the summaries $t$ instead of the data $y$ unless the summaries are lossless. Second, minimizing the EPE implies that the loss functional in Eq. 2 is the expected KL divergence. Similar to the MI, evaluating the expected KL divergence is challenging and neither $f\left(\theta\mid z\right)$ nor $f\left(\theta\mid t\right)$ are known in practice.
To summarize, minimizing the EPE, maximizing the MI between parameters $\theta$ and summaries $t$, maximizing the expected surprise, and minimizing the expected KL divergence between $f\left(\theta\mid z\right)$ and $f\left(\theta\mid t\right)$ are equivalent. But minimizing the EPE is preferable because it can be estimated using Eq. 4 for functional approximations of the posterior and nearest-neighbour entropy estimators for posterior samples (Singh et al., 2003).

### v4

@@ Abstract
Extracting low-dimensional summary statistics from large datasets is essential for efficient (likelihood-free) inference. We characterize three different classes of summaries and demonstrate their importance for correctly analyzing dimensionality reduction algorithms. We demonstrate that minimizing the expected posterior entropy (EPE) under the prior predictive distribution of the model provides a unifying principle that subsumes many existing methods; they are shown to be equivalent to, or special or limiting cases of, minimizing the EPE. We offer a unifying framework for obtaining informative summaries and propose a practical method using conditional density estimation to learn high-fidelity summaries automatically. We evaluate this approach on diverse problems, including a challenging benchmark model with a multi-modal posterior, a population genetics model, and a dynamic network model of growing trees. The results show that EPE-minimizing summaries can lead to posterior inference that is competitive with, and in some cases superior to, dedicated likelihood-based approaches, providing a powerful and general tool for practitioners.
@@ 1 Introduction
Empowered by advances in both scientific understanding and computing, researchers are developing ever more sophisticated simulators. For example, simulated weak lensing maps capture how dark matter affects light propagating through the universe [Merten2019, Fluri2021], coalescent simulators predict the evolution of genetic material [Nordborg2019Coalescent], and synthetic networks shed light on political opinion formation [Sobkowicz2012], effective vaccination strategies [Yang2019], and interactions between proteins [Grassmann2024ProteinInteractions].
While simulators can generate data $y$ given parameters $\theta$, we are often interested in the inverse problem: Constraining parameters $\theta$ given data $y$. If the likelihood $g\left(y\mid\theta\right)$ is available, we can use Markov chain Monte Carlo samplers [Carpenter2017] or variational inference [Bishop2006, Ch. 10] to investigate the posterior $f\left(\theta\mid y\right)$. But inference is more challenging if the likelihood is intractable or costly to evaluate.
Approximate Bayesian computation (ABC) overcomes this challenge in three steps by comparing observed with simulated data [Beaumont2019]: First, we draw many samples $\left(\theta_{i},z_{i}\right)$ from the prior predictive distribution which form the so-called reference table. Second, we evaluate the distance $d_{i}=d\left(y,z_{i}\right)$ between observed data $y$ and the $i^{\mathrm{th}}$ simulated dataset $z_{i}$. Finally, we accept $\theta_{i}$ as a sample from the ABC posterior $\tilde{f}\left(\theta\mid y\right)$ if the distance $d_{i}$ is smaller than a threshold $\epsilon$. The smaller $\epsilon$, the better the approximation. Intuitively, ABC samples parameters $\theta_{i}$ that generate data $z_{i}$ which “look like” the observed data $y$. Hereafter, $y$ and $z$ will denote observed and simulated data.
@@
Figure 1: Different methods for compressing data to informative summaries are intimately related; distinguishing between classes of summaries is essential. Panel (a) illustrates that five information-theoretic approaches (ITAs) are equivalent. They implicitly minimize the same loss (Sections 2 and 3). Approximate sufficiency (Section 4.1) seeks to achieve lossless compression, and minimizing the posterior entropy (Section 4.2) is a special case of ITAs focused on only the observed data. Maximizing Fisher information (Section 4.3) and minimizing $L^{2}$ Bayes risk (Section 4.4) are equivalent each other and ITAs in the large-sample limit. Probabilistic model selection (Section 4.6) maps onto ITAs if we treat model labels as parameters. A dashed arrow from one method to another indicates that the latter is a specialization of the former. Solid arrows indicate correspondence in the large-sample limit. Panel (b) illustrates relationships between classes of summaries. Sufficient statistics $\mathcal{S}$ are a subset of lossless statistics $\mathcal{L}$ although the former only exist if the likelihood belongs to the exponential family. The intersection of lossless summaries $\mathcal{L}$ and the summaries $\mathcal{T}$ considered by the practitioner are optimal summaries $\mathcal{O}$. Optimal summaries are not necessarily lossless, e.g. if $\mathcal{T}$ is restricted to parametric transformations.
@@
A plethora of methods has been developed to address this question; some are summarized in panel (a) of Fig. 1. They include methods to select informative summaries from a pool of candidates [Blum2010, Joyce2008, Nunes2010, Barnes2012, Blum2013] and parameterized transformations that can be optimized to learn summaries [Aeschbacher2012, Fearnhead2012, Prangle2014, Jiang2017, Chan2018, Charnock2018, Chen2021MutualInfo, Radev2022BayesFlow]. Loss functionals quantifying how well the compressor preserves information have been motivated by minimizing the Bayes risk [Fearnhead2012, Jiang2017], model selection [Prangle2014, Raynal2023, Merten2019], and information theoretic arguments [Nunes2010, Chen2021MutualInfo, Barnes2012, Charnock2018, Radev2022BayesFlow].
@@
We characterize three different classes summaries in Section 2: sufficient, lossless, and optimal summaries. In Section 3, we argue that all information-theoretic approaches are equivalent. They implicitly minimize the same loss functional between the summary posterior $f\left(\theta\mid t\right)$ given only $t$ and the true posterior $f\left(\theta\mid y\right)$ given the entire dataset $y$. While these results are well established in information theory, they provide a unifying perspective of different summary extraction approaches. Minimizing the expected posterior entropy (EPE) should be the practitioner’s choice because it is easier to evaluate than either the mutual information (MI) between model parameters and summaries or the Kullback-Leibler (KL) divergence between the posterior given the full data and posterior given only summaries. It also has strong connections with conditional posterior density estimation [Papamakarios2016, Lueckmann2017]. But even methods developed to address different problems (such as parameter inference or model selection) in diverse fields (such as cosmology or population genetics), have strong ties to information-theoretic approaches. For example, in Section 4 we show that maximizing the determinant of the Fisher information [Heavens2000, Charnock2018] and minimizing the $L^{2}$ Bayes risk [Fearnhead2012, Jiang2017] are both equivalent to minimizing the EPE in the large-sample limit. Similarly, learning a probabilistic classifier for model selection [Prangle2014] minimizes the EPE. In Section 5, we discuss concrete steps for learning summaries by fitting conditional posterior density estimators to simulated data. To compare different methods, we devise a benchmark problem with simple likelihood but data that prove challenging for summary selection in Section 5.2. We also compare summary selection approaches on two applied examples: Inferring the mutation and recombination rates of a population genetics model (Section 5.3) and the attachment kernel for a model of growing trees (Section 5.4).
@@ 2 Background
Given data $y$ we seek to infer parameters $\theta$ of a model using summaries $t=t(y)$ that retain as much information about the true posterior as possible. Summaries $t_{\text{suff}}$ with fixed and finite dimensions are Bayes sufficient if $f\left(\theta\mid t_{\text{suff}}\right)=f\left(\theta\mid y\right)$ for all $y$ and any prior $\pi\left(\theta\right)$ [Prangle2018]. But they only exist for exponential-family likelihoods [Koopman1936]. We have to relax the concept of sufficiency, and we call statistics $t_{\text{lossless}}$ lossless if
@@
for all data $y$ of the same sample size, and a given prior $\pi\left(\theta\right)$. While lossless statistics always exist (e.g. the identity map), they may not be useful in practice. We say that the statistics $t_{\text{opt}}$ are optimal if they minimize a non-negative loss functional that measures the discrepancy between the posterior given the full data and the posterior given only summaries. Specifically, we consider the loss functional
$\mathcal{L}_{t}=\int dz\,q\left(z\right)\ell\left\{f\left(\theta\mid z\right),f\left(\theta\mid t(z)\right)\right\},$ (2)
where $\ell$ is an instance-level loss functional that measures the discrepancy between true posterior $f\left(\theta\mid z\right)$ and summary posterior $f\left(\theta\mid t(z)\right)$ for a particular dataset $z$. Instance-level discrepancy measures $\ell$ include, for example, the KL divergence, Wasserstein distance, and total variation distance [Cai2022]. As we discuss further in Section 4.5, summaries that are informative for one dataset may be uninformative for another. The weighting function $q$ encodes which parts of the data space we prioritize. The optimal summaries are
$t_{\text{opt}}=\operatorname{argmin}_{t\in\mathcal{T}}\mathcal{L}_{t},$ (3)
@@
where $\mathcal{T}$ is the space of summaries under consideration. Consequently, sufficient statistics are lossless, and lossless statistics are optimal, but the converse is not necessarily true. For example, $\mathcal{T}$ may be restricted to parametric transformations [Fearnhead2012] or selecting at most $k$ summaries from a set of candidate statistics [Raynal2023]. The relationship between different classes of summaries is illustrated in panel (b) of Fig. 1.
The choice of summary statistic $t$ imposes a fundamental limit on the fidelity of the resulting posterior approximation irrespective of the ABC tolerance $\epsilon$. In the limit $\epsilon\rightarrow 0$, the distribution of accepted samples converges to the summary posterior $f\left(\theta\mid t(y)\right)$. This distribution represents the best possible posterior approximation achievable with a given set of summaries. Consequently, even an ideal ABC procedure cannot recover information about the parameters that is lost during the initial data compression step. Minimizing the loss functional in Eq. 3 improves this asymptotic target, ensuring that the best-case outcome of the inference is a high-fidelity approximation of the true posterior $f\left(\theta\mid y\right)$.
@@
Despite the pursuit of the holy grail of sufficient statistics, we typically have to settle for the weakest concept of optimal statistics. Even the most sophisticated method cannot extract sufficient statistics if the likelihood does not belong to the exponential family [Koopman1936]. Similarly, unless the family of summaries $\mathcal{T}$ is rich enough, lossless compression is not achievable. Further, even if $\mathcal{T}$ is rich enough, one cannot in general verify that Eq. 1 holds for all $\theta$ and $y$ given a finite computational budget.
@@
While models with exponential-family likelihoods are theoretically appealing, they may not be sufficiently expressive or intuitive to address real-world problems. Domain knowledge can aid in the development of models that capture salient features of the data, including protein interaction networks [Grassmann2024ProteinInteractions], cosmology [Charnock2018], and population-genetics [Nordborg2019Coalescent]. But these models often do not have sufficient statistics or even tractable likelihoods, and we need to resort to possibly lossy compression and likelihood-free inference.
@@ 3 Minimizing the expected posterior entropy
A natural loss functional to minimize is the expected KL divergence from the true posterior $f\left(\theta\mid z\right)$ to the summary posterior $f\left(\theta\mid t(z)\right)$. Similar to the evaluation of the Fisher information [Bishop2006, Ch. 6], the expectation is taken with respect to the prior predictive distribution $p(z)$ of the model, i.e.q(z)=p(z)q(z)=p(z). This ensures that the summaries are informative for data that are plausible under the model. We propose choosing summaries that minimize the expected posterior entropy (EPE). This approach is equivalent to minimizing the expected KL divergence, conceptually simple, computationally tractable, and has a strong connection with recent inference techniques based on conditional density estimation [Papamakarios2016, Lueckmann2017, Radev2022BayesFlow].
@@
Here, a fiducial dataset refers to a dataset generated based on known parameters. Taking the expectation with respect to the data under the model yields the EPE
@@
where the Jacobian has been absorbed by the joint density $p\left(t,\theta\right)$. With a slight abuse of notation, we use $p\left(\cdot\right)$ for both the marginal likelihood and joint distribution where the distinction is unambiguous. Given a posterior density estimator $\hat{f}\left(\theta\mid t\right)$ that seeks to approximate the summary posterior, we can construct a Monte Carlo estimate of the EPE
$\hat{\mathcal{H}}{}=-m^{-1}\sum_{i=1}^{m}\log\hat{f}\left(\theta_{i}\mid t(z_{i})\right),$ (5)
where $\theta_{i}$ and $z_{i}$ are joint samples from $p\left(\theta,z\right)$, and $m$ is the number of samples. This estimate is the widely used loss function for learning the posterior from simulated data [Papamakarios2016, Lueckmann2017, Radev2022BayesFlow], where $m$ is the size of the mini-batch, i.e. a subset of the data used to train the model.
We consider three well-established connections to other information-theoretic approaches [Bishop2006, Ch. 1] although with a specific focus on the selection of summaries for ABC. First, we evaluate the difference between the prior entropy and EPE
@@
by the law of total probability. The inner integral of Eq. 6 is the KL divergence from the prior to the posterior $D_{\mathrm{KL}}\left(f\left(\theta\mid t\right)\;\|\;\pi\left(\theta\right)\right)$, sometimes called surprise because it measures the degree to which an observer updates their belief in light of new data [Itti2009]. Minimizing the EPE thus maximizes our expected surprise from observing the summaries because the prior entropy does not depend on the choice of summaries.
@@
As the MI is non-negative, the EPE is not larger than the prior entropy, i.e. we reduce uncertainty on average. Minimizing the EPE is equivalent to maximizing the MI which has been proposed in the context of subset selection [Barnes2012] and neural summaries [Chen2021MutualInfo]. However, estimating MI is difficult in high dimensions [Jeffrey2020NeuralSummaries], making the approach computationally challenging.
@@
and we can identify the inner integral as the KL divergence from the summary posterior $f\left(\theta\mid t\right)$ to the true posterior $f\left(\theta\mid z\right)$ (see App. A). The difference of expected entropies is thus equal to the expected KL divergence between the posteriors
@@
which Chan2018 used to infer recombination hotspots in population genetics and Radev2022BayesFlow targeted for amortized Bayesian inference. Minimizing the EPE is equivalent to minimizing the expected KL divergence because the true posterior entropy given the complete dataset does not depend on the summaries. The KL divergence is non-negative which allows us to draw two conclusions. First, the EPE given only summaries $t\left(z\right)$ is greater than or equal to the EPE given the full dataset $z$, i.e. we lose information in expectation by conditioning on the summaries $t$ instead of the data $y$ unless the summaries are lossless. Second, minimizing the EPE implies that the loss functional in Eq. 3 is the expected KL divergence. Similar to the MI, evaluating the expected KL divergence is challenging because neither the true posterior $f\left(\theta\mid z\right)$ nor the summary posterior $f\left(\theta\mid t\left(z\right)\right)$ are known in practice.
To summarize, minimizing the EPE, maximizing the MI between parameters $\theta$ and summaries $t$, maximizing the expected surprise, and minimizing the expected KL divergence between $f\left(\theta\mid z\right)$ and $f\left(\theta\mid t\left(z\right)\right)$ are equivalent, as illustrated in panel (a) of Fig. 1. But minimizing the EPE is preferable because it can be estimated using Eq. 5 for functional approximations of the posterior and nearest-neighbor entropy estimators for posterior samples [Singh2003].